OpenStudy (ray10):

use partial derivatives to approximate the value of \[f(6.1, 1.9)\] where \[f(x,y) = (x^{2}-y^{3}-1)^{\frac{1}{3}}\] at \[(6,2)\]

terenzreignz (terenzreignz):

I wonder how much longer I should stare at this... ^_^

terenzreignz (terenzreignz):

Well, @Ray10 ... ready when you are:

OpenStudy (ray10):

oh sorry!!! My laptop froze and I didn't realize!! :O @terenzreignz

OpenStudy (ray10):

I am ready :D

terenzreignz (terenzreignz):

Well I'm not. You took a while. Just let me patch things up with her http://openstudy.com/study#/updates/52319b0de4b03eb771a27ef9 (don't worry, it's just algebra) and I'll get back to you. I'll do my best to hurry it along, but remember, her learning is a priority ^_^

OpenStudy (ray10):

yes sure thing TJ! :D of course! I understand :) I'll be waiting

OpenStudy (ray10):

you sent me the link to my own question :P

terenzreignz (terenzreignz):

Whoops... well if you're curious, check this http://openstudy.com/study#/updates/5231a34de4b012524e87f0c4

OpenStudy (ray10):

you teach very well :)

terenzreignz (terenzreignz):

Thank you ^_^

terenzreignz (terenzreignz):

You know what? While you're at it, find \(\Large f_x\) and \(\Large f_y\) They'll be needed.

terenzreignz (terenzreignz):

Okay, I'm basically done. Have you found these partials?

OpenStudy (ray10):

so far this is what I have got: \[f(x,y) = (6,2) => f(x,y)=3\] I found fx and fy to be \[fx=\frac{ 2x }{( \sqrt[3]{3(x^{2}-y^{3}-1)} })^{2}\] and \[fy=\frac{ -y^{2} }{(\sqrt[3]{x^{2}-y^{3}-1} )^{2}}\] \[fx=(x=6), = \frac{ 4 }{ 9 }\] so far correct ?

OpenStudy (ray10):

small typo at fx, the bracket is meant to be on the denominator :P

terenzreignz (terenzreignz):

I hate fractional exponents -.-

OpenStudy (ray10):

oh me too!! -.-

terenzreignz (terenzreignz):

By the way... f_x for subscripts. \[\Large f(x,y) = (x^{2}-y^{3}-1)^{\frac{1}{3}}\]

terenzreignz (terenzreignz):

\[\Large f_x(x,y)=\frac13(x^2-y^3-1)^{-\frac23}\cdot2x\] right?

OpenStudy (ray10):

yes that is correct :)

terenzreignz (terenzreignz):

Find \(\large f_y\) And for heaven's sake, don't use radicals.

OpenStudy (ray10):

what are radicals? :S

terenzreignz (terenzreignz):

These nasty things: \[\Huge \color{red}{\sqrt{\qquad}}\]

OpenStudy (ray10):

\[fy(x,y)=1/3(x^{2}−y^{3}−1)^{−2/3}⋅-y^{2}\]

terenzreignz (terenzreignz):

Fraction \frac{numerator}{denominator} subscripts f_x etc

OpenStudy (ray10):

why is your formula's nice and big to read, and mine are small? :S

terenzreignz (terenzreignz):

increase size... \large \Large \LARGE \huge \Huge before the whole thing.

terenzreignz (terenzreignz):

I use \Large by default

terenzreignz (terenzreignz):

Okay, good. Enter: the total differential: \[\Large dz = f_x(x_o,y_o)\partial x + f_y(x_o,y_o)\partial y\]

OpenStudy (ray10):

\[\Large \frac{ 1 }{ 3 }(x^{2}-y^{3}-1)^\frac{ 2 }{ 3 } \times -y^{2}\] ^_^

terenzreignz (terenzreignz):

Here, of course, the partial derivatives are evaluated at the point (6,2) What are \(\large f_x(6,2)\) and \(\large f_y(6,2)\)?

OpenStudy (ray10):

When I use that total differential, I see to arrive at: \[\Large f(6.1,1.9)=f(6,2)+(0.1)\frac{ 4 }{ 9 }+(-0.1)\frac{ 4 }{ 9 }\]

terenzreignz (terenzreignz):

You get 4/9 at both partial derivatives?

OpenStudy (ray10):

For \[\Large fx\] I get \[\Large \frac{ 4 }{ 9 }\] and for \[\Large fy\] I get \[\Large -\frac{ 4 }{ 9 }\] is that correct?

terenzreignz (terenzreignz):

abusing the size option? :P

OpenStudy (ray10):

haha sorry, it looks neater :P I'll dial it down :P

terenzreignz (terenzreignz):

You might want to redo your \(\Large f_y\) <again and again... I told you that to make subscripts, you just have f_{subscript}>

OpenStudy (ray10):

\[\large f_{y}= \frac{ 1 }{ 3 }(x^{2}-y^{3}-1)^\frac{ -2 }{ 3 } \times -y^{2}\] \[\large f_{y}= \frac{ 1 }{ 3 }(6^{2}-2^{3}-1)^\frac{ -2 }{ 3 } \times -2^{2}\] \[\large f_{y} = \frac{ 4 }{ 27 }\] is what I get

terenzreignz (terenzreignz):

I wonder how the negative sign... just disappeared... strange :3 \[\large f_{y}= \frac{ 1 }{ 3 }(6^{2}-2^{3}-1)^\frac{ -2 }{ 3 } \times \boxed-2^{2}\]

terenzreignz (terenzreignz):

Carelessness is your worst enemy. ^_^

OpenStudy (ray10):

ah I keep forgetting to type it in -.- ah it's a task to type each line out :P

terenzreignz (terenzreignz):

I wonder how the negative sign... just disappeared... strange :3 \[\large f_{y}= \frac{ 1 }{ 3 }(6^{2}-2^{3}-1)^\frac{ -2 }{ 3 } \times \boxed-2^{2}\]

OpenStudy (ray10):

\[\large \frac{-4}{27}\] correct? :D

OpenStudy (ray10):

I then still don't get the correct final answer :(

terenzreignz (terenzreignz):

What is it?

terenzreignz (terenzreignz):

I mean, the correct final answer

OpenStudy (ray10):

the answer should be around 6.39 :/ I get 3....

terenzreignz (terenzreignz):

Impossible.

OpenStudy (ray10):

to work out the correct answer, it's just: \[\large (6.1^{2}+1.9^{2})^\frac{ 1 }{ 2 }\] ??

terenzreignz (terenzreignz):

<smh> The correct answer is \[\Large f(6.1,1.9)=(6.1^2-1.9^3-1)^{\frac13}\] How on earth did you get\[\large (6.1^{2}+1.9^{2})^\frac{ 1 }{ 2 }\]???

OpenStudy (ray10):

ohhhhhh there was a misunderstanding in my lecture notes!! let me show you :P

terenzreignz (terenzreignz):

Yes... show me~ good idea :)

OpenStudy (ray10):

I thought the bottom line is a standard equation to work out the actual value -.-

OpenStudy (ray10):

this is yet another good reason for not doing work soo late into the night :P

terenzreignz (terenzreignz):

You do realise that this is an entirely different function, right? -.-

OpenStudy (ray10):

oh yeah of course! That is just the lecture notes. I was applying it to another question and was following the method :S

terenzreignz (terenzreignz):

Oh... too methodical :P I haven't taken notes (in maths) since eighth grade :P

terenzreignz (terenzreignz):

Anyway, do the total differential again. You were close.

OpenStudy (ray10):

\[\Large f(6.1,1.9)=f(6,2)+(0.1)\frac{ 4 }{ 9 }+(-0.1)\frac{ -4 }{ 27 } = 3.06\] \[\Large :D\]

terenzreignz (terenzreignz):

Aren't you glad this all worked out in the end? :)

OpenStudy (ray10):

I am more than glad, you seem to always give that extra external insight which helps the poster, work out the problem :D

terenzreignz (terenzreignz):

By the way, allow me to nitpick...\[\Large f(6.1,1.9)\color{red}\approx f(6,2)+(0.1)\frac{ 4 }{ 9 }+(-0.1)\frac{ -4 }{ 27 } \color{red}\approx 3.06\]

terenzreignz (terenzreignz):

Reminding you that this is JUST an approximation :P

OpenStudy (ray10):

Oh of course!! :P Thank you for reminding! And thank you very much foryour help!! :D

terenzreignz (terenzreignz):

No problem ^_^

terenzreignz (terenzreignz):

Here's that insight. If something changes in x and y, something will change in z = f(x,y) \[\Large f(x+\Delta x , y+\Delta y) =f(x,y)+\color{red}{\Delta z}\]

terenzreignz (terenzreignz):

Now derivatives are all about differences, and \(\large \Delta z\) is approximated by \[\Large dz = f_x\partial x + f_y \partial y\]

terenzreignz (terenzreignz):

Which in turn is approximated by \[\Large dz \approx f_x\Delta x + f_y \Delta y \]

terenzreignz (terenzreignz):

And that's where THIS approximation comes from: \[\Large f(x+\Delta x , y+\Delta y) \color{green}\approx f(x,y)+\color{red}{f_x(x,y)\Delta x + f_y(x,y) \Delta y }\]

OpenStudy (ray10):

ahhh there we go!! Now I have a clearer understanding :D thank you TJ!! sorry for the late reply!! My computer crashed, so I had to restart it and fix it up :/ @terenzreignz